OpenStudy (anonymous):

If y varies inversely as the square of x, and y = 1/8 when x = 1, find y when x = 5.

OpenStudy (anonymous):

so the formula is y=k/x

OpenStudy (anonymous):

so 1/8=k/(1^2)

OpenStudy (anonymous):

do you understand what im saying?

OpenStudy (anonymous):

@Timtime

OpenStudy (anonymous):

y = 1/200 y = 1/125 y = 1/5 y = 8/25

OpenStudy (anonymous):

those are the multiple choice answers iam stuck

OpenStudy (anonymous):

ok what is 1^2?

OpenStudy (anonymous):

how u get 1^2

OpenStudy (anonymous):

square of x x=1

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

so what is it?

OpenStudy (anonymous):

1^2? that would just equal 1

OpenStudy (anonymous):

correct, so multiply both sides by 1 to get k

OpenStudy (anonymous):

so k=1/8, now enter in the new info. y=(1/8)/5 and solve fir y

OpenStudy (anonymous):

for

OpenStudy (anonymous):

is it 8/25

OpenStudy (anonymous):

(1/8)/25 or 1/200

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

not a problem